OpenStudy (anonymous):

tan x=1/2a, pi <= x <= 3pi/2, determine an exact value for cos(2x)

OpenStudy (mathmath333):

is it this \(tanx=\dfrac{1}{2a}\) or this \(tanx=\dfrac{1}{2}a\)

OpenStudy (anonymous):

The first, sorry I should have added brackets.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{tan x=\dfrac{1}{2a}\\~\\ tan^2x=\dfrac{1}{4a^2}\\~\\ tan^2x+1=\dfrac{1}{4a^2}+1\\~\\ sec^2x=\dfrac{1+4a^2}{4a^2}\\~\\ cos^2x=\dfrac{4a^2}{1+4a^2}----- \color{red}{1}\\~\\ 1-sin^2x=\dfrac{4a^2}{1+4a^2}\\~\\ -sin^2x=\dfrac{4a^2}{1+4a^2}-1----- \color{red}{2}\\~\\ \text{by adding 1 and 2}\\~\\ cos^2x-sin^2x=\dfrac{4a^2}{1+4a^2}+\dfrac{4a^2}{1+4a^2}-1\\~\\ cos(2x)=\dfrac{4a^2}{1+4a^2}+\dfrac{4a^2}{1+4a^2}-1\\~\\ =\dfrac{8a^2-1-4a^2}{1+4a^2}\\~\\ =\dfrac{4a^2-1}{4a^2+1}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Ahh okay, I see what you did. I was going about it the wrong way. Thank you kindly. It's much appreciated.

OpenStudy (mathmath333):

yw